how many ml of 0.1m hcl is requied to react completly with 1gr mixture of na2co3 and nahco3 contining equi molar amount of both?
 

Grade:11

1 Answers

Vikas TU
7 years ago
let x gm of Na2CO3 is present then 1-x is remaining NaHCO3.
And equal moles are given therfore,
x/ 106 = (1-x)/ 84
on solving we get,
x = 0.558 gm.
Hence on putting in the moles rspectively we get,
moles = 0.0152 for each.
And the number of moles fo HCl required would be => 2*0.0512 => 0.01502
Total moles would be => 0.01502 + 0.00525 => 0.015728 moles.
Hence,
MV = moles we get,
V = 0.015728/0.1 => 0.15728 L.